40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 19 16:00:18 2010, Jenne, Update, IOO, Hooray! We locked the MC! (and some other stuff) 
    Reply  Wed May 19 22:28:05 2010, Koji, Update, IOO, Hooray! We locked the MC! (and some other stuff) 
Message ID: 2952     Entry time: Wed May 19 16:00:18 2010     Reply to this: 2954
Author: Jenne 
Type: Update 
Category: IOO 
Subject: Hooray! We locked the MC! (and some other stuff) 

[Jenne, Kevin]

We opened up the MC chambers again, and successfully got the MC locked today!  Hooray!  This meant that we could start doing other stuff....

First, we clamped the Faraday.  I used the dog clamps that Zach left wrapped in foil on the clean cart.  I checked with a card, and we were still getting the 00 mode through, and I couldn't see any clipping.  2 thumbs up to that.

Then we removed the weight that was on the OMC table, in the way of where MMT2 needs to go.  We checked the alignment of the MC, and it still locks on TEM00, but the spot looks pretty high on MC2 (looking at the TV view). We're going to have to relevel the table when we've got the MMT2 optic in the correct place.

We were going to start moving the PZT steering mirror from the BS table to the IOO table, place MMT2 on the OMC table, and put in a flat mirror on the BS table to get the beam out to the BS oplev table, but Steve kicked us out of the chambers because the particle count got crazy high.  It was ~25,000 which is way too high to be working in the chambers (according to Steve).  So we closed up for the day, and we'll carry on tomorrow. 

 

Photos of the weight before we removed it from the OMC table, and a few pictures of the PZT connectors are on Picasa

ELOG V3.1.3-